# A Practitioner’s Guide to Kolmogorov–Arnold Networks

### v2

@@ 3 KST Toward Neural Networks
Approximate KST (inner–and–outer; controversial). These approaches approximate or entirely replace the universal inner functions, thereby departing from the exact Kolmogorov formulation.
  • Kurková (1991–1992): replaces the universal inner staircase by sigmoidal staircase approximants and learns the outer coefficients within a two-hidden-layer sigmoidal architecture [90, 91].
  • Nees (1994, 1996): constructs explicit recursive piecewise-linear approximations of the inner maps $\phi_{q}^{k}$ together with data-driven outer corrections $g_{r}$ that achieve geometric error decay [102, 103].
  • Spline-based architectures: The KSN model of Igelnik and Parikh (2003) [106] uses cubic splines for both inner and outer functions, while Coppejans (2005) [107] introduces monotone inner splines for statistical identifiability. Related spline-based superpositional models, including ExSpliNet [108], KASAM [109], and the Urysohn-operator framework of Polar and Poluektov [110], adopt additive or hierarchical structures reminiscent of KST but do not enforce universality of the inner layer. These methods yield practical spline architectures but are not constructive realizations of the Kolmogorov superposition theorem.
@@
Neural interpretations. De Figueiredo (1980) [104] was the first to view the Kolmogorov–Sprecher representation as a computational architecture, expressing it in block-diagram form and proposing a data-driven procedure for learning the outer function $g$ using Chebyshev bases. Hecht–Nielsen (1987) [105] interprets Sprecher’s representation as a three-stage mapping network with a fixed universal hidden layer and learned outer functions. Kurková’s approximation framework (1991–1992) [90, 91] provides compatibility with standard trainable neural networks by replacing the inner staircase with smooth sigmoidal combinations while learning the outer layer. Spline-based models such as KSN (2003) [106] and the shape-constrained variant of Coppejans (2005) [107] train both inner and outer functions within spline families, yielding practical but non-KST architectures. See Table 4 for a chronological overview of these developments, tracing the progression from the original KST theory to its computational and neural interpretations.
@@
Liu et al. [26] introduce the Kolmogorov–Arnold Network, a neural architecture explicitly motivated by the structure of the KST (4), though not intended as a mathematical realization of it. The authors emphasize that KAN is not an exact implementation of the Kolmogorov Superposition Theorem, but rather a computational architecture “inspired by” the Kolmogorov–Arnold representation.
@@ 4 Bridging KANs and MLPs
Formal Equivalence. Wang et al. [53] establish a bidirectional correspondence between MLPs and KANs. Any MLP with $\mathrm{ReLU}^{k}$ activations,
$\sigma_{k}(x)=\max(0,x)^{k},$
can be represented by an equivalent KAN parameterization using B-spline bases of order $k$ and grid size G=2G{=}2. For an MLP approximating a target function $f:\mathbb{R}^{d}\!\to\!\mathbb{R}$,
[…]
there exists an equivalent KAN representation,
[…]
where each univariate function $\varphi_{ij}:\mathbb{R}\!\to\!\mathbb{R}$ is a learned B-spline that approximates $\sigma_{k}(w_{ij}x_{j}+b_{i})$. In essence, KANs replace the fixed nonlinearities of MLPs with learnable one-dimensional functions, enabling richer local adaptation without increasing architectural depth.
The reverse direction also holds under mild constraints. If a KAN layer employs polynomial-type univariate functions (e.g., B-splines of order $k$) and excludes non-polynomial activations such as SiLU, the model can be reformulated as an equivalent MLP with $\mathrm{ReLU}^{k}$ activations. In this conversion, a KAN of width $W$, grid size $G$, and depth $L$ induces an MLP whose effective width grows as $(G+2k+1)W^{2}$, resulting in a parameter complexity of $\mathcal{O}(G^{2}W^{4}L)$. By comparison, the original KAN requires only $\mathcal{O}(GW^{2}L)$ parameters—highlighting its superior parameter efficiency, especially for fine grids or high-order spline bases. This equivalence creates a bridge through which approximation and convergence results from MLP theory can be transferred directly to KANs (see Section 10).
@@
5 How KANs Extend the MLP Landscape
@@ 6.1 B-spline
Subfigures 1(a) and 1(b) visualize cubic (k=3k{=}3) B-spline bases on non–extended and extended grids, respectively. Without extension, boundary bases are truncated and lose symmetry; with extension, added knots outside the domain restore full polynomial support (gray bands). To compare expressivity, subfigures 1(c) and 1(d) synthesize univariate maps
@@
Spectral bias and NTK conditioning. Spline grids control frequency reach: finer grids reduce spectral bias but may worsen conditioning [53, 39]. Initialization also matters—Rigas et al. [276] [^14] https://github.com/srigas/KAN_Initialization_Schemes show that Glorot-style and power-law initializations yield more stable NTKs and faster early convergence. (See Secs. 10.2 and 10.3.)
@@ 6.3 ReLU
Figure 3: (a) Step-by-step construction of the normalized ReLU-based local basis function from its constituent $\operatorname{ReLU}(e_{i}-x)$ and $\operatorname{ReLU}(x-s_{i})$ terms. (b) Complete set of normalized ReLU–KAN basis functions $R_{i}(x)$ with supports $s_{i},e_{i}] [$123].

### v3

@@ 2 Bridging KANs and KST
Approximate KST (inner–and–outer). These approaches approximate or replace the universal inner functions of the Kolmogorov superposition theorem, thereby departing from the exact Kolmogorov formulation while enabling practical approximation and learning.
  • Kurková (1991–1992): initiated the approximation-theoretic study of the Kolmogorov superposition theorem by interpreting Kolmogorov’s structure in the sense of approximation rather than exact representation. In the Kolmogorov–Sprecher form, the universal inner functions need only satisfy prescribed values on finitely many intervals and can therefore be approximated by staircase-like combinations of a sigmoidal activation. This yields a fully sigmoidal analogue of the Kolmogorov scheme and resolves earlier concerns regarding non-smooth inner maps and non-parametric outer functions [94, 95].
    Theorem 4 (Kurková 1992 [95])
    Let $n\geq 2$, let $\sigma:\mathbb{R}\to I$ be sigmoidal, and let $f\in C(I^{n})$. For every $\varepsilon>0$ there exist $k\in\mathbb{N}$ and functions $\varphi_{i},\psi_{p,i}\in S(\sigma)$ such that
    $\Bigl|f(x_{1},\ldots,x_{n})-\sum_{i=1}^{k}\varphi_{i}\!\Bigl(\sum_{p=1}^{n}\psi_{p,i}(x_{p})\Bigr)\Bigr|<\varepsilon,\qquad\forall(x_{1},\ldots,x_{n})\in I^{n}.$
    Thus, the Kolmogorov superposition $\sum_{p}\phi_{p,q}(x_{p})$ can be replaced by smooth sigmoidal compositions, yielding learnable inner and outer functions within a two-hidden-layer architecture. Kurková’s results provided the first rigorous bridge between Kolmogorov-type superpositions and standard neural network approximation, demonstrating both theoretical universality and practical realizability.
  • Nees (1994, 1996): constructed explicit recursive piecewise-linear approximations of the inner maps $\phi_{q}^{k}$ together with data-driven outer corrections $g_{r}$, achieving geometric error decay [106, 107].
  • Spline-based architectures: The KSN model of Igelnik and Parikh (2003) [110] uses cubic splines for both inner and outer functions, while Coppejans (2005) [111] introduces monotone inner splines for statistical identifiability. Related spline-based superpositional models, including ExSpliNet [112, 113], KASAM [114], and the Urysohn-operator framework of Polar and Poluektov [115], adopt additive or hierarchical structures reminiscent of KST but do not enforce universality of the inner layer.
@@
Neural interpretations. De Figueiredo (1980) [108] was the first to view the Kolmogorov–Sprecher representation as a computational architecture, expressing it in block-diagram form and proposing a data-driven procedure for learning the outer function $g$ using Chebyshev bases. Later, Hecht–Nielsen (1987) [109] interprets Sprecher’s representation as a three-stage mapping network with a fixed universal hidden layer and learned outer functions. Kurková’s approximation framework (1991–1992) [94, 95] provides compatibility with standard trainable neural networks by replacing the inner staircase with smooth sigmoidal combinations while learning the outer layer. Spline-based models such as KSN (2003) [110] and the shape-constrained variant of Coppejans (2005) [111] train both inner and outer functions within spline families, yielding practical but non-KST architectures. See Table 3 for a chronological overview of these developments, tracing the progression from the original KST theory to its computational and neural interpretations.
@@
Liu et al. [26, 27, 28] introduce the Kolmogorov–Arnold Network, a neural architecture explicitly motivated by the structure of the KST (1), though not intended as a mathematical realization of it. The authors emphasize that KAN is not an exact implementation of the Kolmogorov Superposition Theorem, but rather a computational architecture “inspired by” the Kolmogorov–Arnold representation.
@@
3 Bridging KANs and Kernel Methods
Kolmogorov–Arnold networks constructed from explicit basis functions admit a direct and instructive comparison with classical kernel methods. This section establishes a concrete bridge between the two paradigms by focusing on how basis functions are combined, how coefficients are parameterized, and how dimensionality is handled. The comparison shows that a shallow, one-dimensional KAN is mathematically equivalent to a kernel regression model, while deeper or multi-dimensional KANs depart from kernel structure in systematic and consequential ways.
Kernel method perspective. A classical kernel or basis-expansion method approximates a target function $f$ by
$f(x)\;\approx\;\sum_{k}c_{k}\,\psi_{k}(x),$
where $\{\psi_{k}\}$ are fixed basis functions (polynomials, splines, Gaussian RBFs, Fourier modes, etc.) and $c_{k}\in\mathbb{R}$ are scalar expansion coefficients. Given data $\{(x_{n},f_{n})\}_{n=1}^{N}$, one forms the design matrix
$\Psi_{n,k}=\psi_{k}(x_{n}),$
and determines the coefficients $c_{k}$ by solving the linear system
$\Psi\,\mathbf{c}\;\approx\;\mathbf{f},$
where $\mathbf{c}$ and $\mathbf{f}$ denote the vectors of expansion coefficients and sampled function values, respectively. Once the basis is fixed, the problem is linear, convex, and solved directly by linear algebra.
We now compare this structure with KANs in three progressively more complex settings.
Case 1: Output-only KANs for one-dimensional problems coincide with kernel methods.
Consider a KAN applied to a one-dimensional problem, with a single scalar input, no hidden layers, and one output neuron. The network produces an approximation $g$ of the target function $f$, which can be written as
$g(x)=\sum_{k}c_{k}\,\phi_{k}(x),$
where $\{\phi_{k}\}$ denotes the chosen univariate basis functions implemented by the KAN and $\{c_{k}\}$ are trainable scalar coefficients. This representation coincides exactly with the standard form of a kernel or basis-expansion model.
The interpretation depends only on the selected basis:
  • for spline-based KANs, $\phi_{k}$ are spline basis functions and $c_{k}$ are spline coefficients;
  • for polynomial-based KANs (e.g. Chebyshev), $\phi_{k}$ are orthogonal polynomials and $c_{k}$ are polynomial coefficients;
  • for Gaussian-based KANs, $\phi_{k}$ are fixed radial basis functions and $c_{k}$ are their amplitudes.
In this output-only setting, the KAN is linear in its parameters and spans the same function class as a classical kernel or basis-expansion model. The coefficient vector $\mathbf{c}$, however, is obtained differently: KANs learn $\mathbf{c}$ via gradient-based optimization of the empirical loss, whereas classical kernel methods typically rely on direct linear-algebraic procedures such as normal equations, matrix inversion, or orthogonal projections. This difference reflects the training strategy rather than the representable function space. As shown in the spectral analysis of shallow KANs in [55], the resulting loss remains quadratic with a Gram-matrix Hessian, enabling conditioning and convergence analysis through linear algebra. Neural-network–specific expressive effects arise only once hidden layers or multi-dimensional compositions are introduced.
Therefore, in the absence of hidden layers, KANs applied to one-dimensional problems are functionally equivalent to classical kernel or basis-expansion models, with differences arising only from the training procedure rather than expressivity.
Case 2: Divergence in multiple dimensions.
The equivalence between KANs and kernel methods no longer holds once the input dimension exceeds one. Classical kernel and basis-expansion methods construct multivariate features by explicitly encoding cross-dimensional interactions. For separable bases, this is achieved through tensor products,
$\psi_{k_{1},\dots,k_{d}}(\mathbf{x})=\prod_{\ell=1}^{d}\psi_{k_{\ell}}(x_{\ell}),$
where $(k_{1},\dots,k_{d})$ is a multi-index and each $\psi_{k_{\ell}}$ is a fixed one-dimensional basis function. For radial basis functions, the dependence is instead on full $d$-dimensional distances,
$\psi_{m}(\mathbf{x})=\varphi\!\big(\|\mathbf{x}-\mathbf{x}_{m}\|\big),\qquad\mathbf{x}_{m}\in\mathbb{R}^{d},$
where $m$ indexes the kernel centers. In both cases, multivariate coupling is embedded directly in the basis functions.
These constructions are expressive but suffer from the curse of dimensionality. For polynomial or spline bases of order $p$, the number of tensor-product basis functions grows as $(p+1)^{d}$, while for radial kernels the number of required centers increases rapidly with $d$ in order to resolve multivariate structure [118, 119].
KANs adopt a different architectural principle. Given an input $\mathbf{x}=(x_{1},\dots,x_{d})$, a standard KAN layer computes the output of neuron $j$ as
$z_{j}=\sum_{\ell=1}^{d}\phi_{j,\ell}(x_{\ell}),\qquad\phi_{j,\ell}\in\mathcal{F}_{1\mathrm{D}},$
where each $\phi_{j,\ell}$ is an independently parameterized univariate function. No tensor-product basis functions $\prod_{\ell}\psi_{k_{\ell}}(x_{\ell})$ and no radial functions $\varphi(\|\mathbf{x}-\mathbf{x}_{m}\|)$ are constructed at the level of a single layer.
Multivariate interactions arise only through the presence of hidden layers. When a second KAN layer is applied to the outputs of the first, the resulting composition produces functions that depend jointly on multiple input coordinates. With increasing depth, such compositions can approximate interaction terms resembling tensor-product or radial features, but these interactions are generated through layer-wise composition rather than encoded explicitly in a single multivariate basis function.
This distinction is structural. As analyzed in [55], the expressive power of KANs in multiple dimensions is governed by depth and compositional structure, whereas classical kernel methods rely on explicit multivariate basis construction. KANs therefore avoid explicit tensorization and achieve expressivity through stacked univariate transformations.
Interpretation. For multivariate problems, KANs can be viewed as structured, low-complexity alternatives to tensor-product or radial kernel spaces. They trade explicit enumeration of all cross-dimensional interactions for a compositional representation that scales more favorably with dimension. This represents a structurally distinct approach to multivariate function approximation rather than a direct kernel approximation.
Case 3: Deep KANs and nonlinear coefficient coupling.
As established in Case 2, hidden layers are required for KANs to represent general multivariate functions. When such hidden layers are introduced, KANs depart most fundamentally from classical kernel and basis-expansion methods. In this deep setting, the coefficients of the resulting function are no longer independent parameters.
This nonlinear coupling arises from function composition across layers and has several consequences:
  • Redundancy. Multiple parameter configurations can represent the same final function, leading to flat directions and equivalent minima.
  • Loss of linear solvability. Unlike kernel methods, the coefficients cannot be recovered by solving a linear system; training requires nonlinear, iterative optimization.
  • Entangled representation. Intermediate features generated by hidden neurons are combined by subsequent layers, obscuring a direct correspondence between individual parameters and the final functional form.
Illustrative example. The effects described above already appear in the simplest one-dimensional deep setting and can be understood through two closely related aspects: growth of the effective polynomial degree and nonlinear coupling of coefficients.
  • Degree growth. Consider a KAN with two Chebyshev layers of degree $P=2$, as in the ChebyKAN construction [136] [^4] https://github.com/SynodicMonth/ChebyKAN, but without any regularization. The first (hidden) layer consists of two neurons, and the second layer produces a scalar output. Although each layer employs only quadratic basis functions, their composition across two layers yields an output polynomial of degree $4$. More generally, for a one-dimensional KAN with $L$ stacked polynomial layers of degree $P$, the effective polynomial degree satisfies
    $\deg(g)=P^{\,L},$
    where $L$ denotes the number of layers. This growth is a direct consequence of function composition and does not arise in shallow models. In this setting, high-degree terms are generated automatically, and training must adjust the parameters so that the corresponding coefficients remain close to zero when the target function does not require such a high polynomial degree.
    However, as illustrated in the item below, the same parameters that control high-degree terms also contribute to lower-degree coefficients, making it nontrivial for optimization to suppress unwanted high-degree components without affecting the lower-degree structure of the solution.
  • Nonlinear coefficient coupling. The same compositional mechanism that increases polynomial degree also changes how the polynomial coefficients depend on the underlying parameters. Let $A^{(1)}\in\mathbb{R}^{3\times 2}$ and $A^{(2)}\in\mathbb{R}^{6\times 1}$ denote the parameter matrices of the first and second layers, respectively, with entries
    $A^{(1)}=\big(a_{ij}\big),\qquad A^{(2)}=\big(b_{ij}\big).$
    For this two-layer Chebyshev–KAN with per-layer polynomial degree $P=2$, collecting powers of $x$ yields the equivalent power-form
    $g(x)=C_{0}+C_{1}x+C_{2}x^{2}+C_{3}x^{3}+C_{4}x^{4},$
    where the coefficients are
    $C_{4}$ $=8{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}}^{2}b_{12}+8a_{22}^{2}b_{15},$
    $C_{3}$ $=8a_{11}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}}\,b_{12}+8a_{21}a_{22}b_{15},$
    $C_{2}$ $=2{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}}\,b_{11}+\big(8a_{10}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}}+2a_{11}^{2}-8{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}}^{2}\big)b_{12}+2a_{22}b_{14}+\big(8a_{20}a_{22}+2a_{21}^{2}-8a_{22}^{2}\big)b_{15},$
    $C_{1}$ $=a_{11}b_{11}+\big(4a_{10}a_{11}-4a_{11}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}}\big)b_{12}+a_{21}b_{14}+\big(4a_{20}a_{21}-4a_{21}a_{22}\big)b_{15},$
    $C_{0}$ $=b_{10}+b_{13}-b_{12}-b_{15}+(a_{10}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}})b_{11}+2(a_{10}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}})^{2}b_{12}+(a_{20}-a_{22})b_{14}+2(a_{20}-a_{22})^{2}b_{15}.$
    For comparison, a classical Chebyshev approximation truncated at degree $4$ has coefficients $\mathbf{c}=(c_{0},\dots,c_{4})^{\top}$ determined by a linear system of the form $\Psi\,\mathbf{c}=\mathbf{f}$ (or its least-squares variant), and admits the power-form
    $f(x)\;\approx\;(c_{0}-c_{2}+c_{4})+(c_{1}-3c_{3})x+(2c_{2}-8c_{4})x^{2}+(4c_{3})x^{3}+(8c_{4})x^{4},$
    where each polynomial coefficient depends linearly on a single entry of $\mathbf{c}$.
    The distinction is therefore explicit. In the classical case, the mapping from data to coefficients is linear and fully described by the design (kernel) matrix $\Psi$. In the deep KAN, each coefficient $C_{k}$ is a nonlinear function of parameters from multiple layers. Terms such as ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}}^{2}$ arise from composition and appear in several coefficients simultaneously, meaning that a single parameter influences multiple polynomial orders. Consequently, increasing depth not only raises the effective polynomial degree, but also couples the coefficients nonlinearly, making their identification a fundamentally different problem from solving a linear kernel system, even in one-dimensional settings.
Summary. Case 1 establishes that shallow KANs applied to one-dimensional problems coincide with classical kernel or basis-expansion models, differing only in training procedure. Case 2 shows that for multivariate inputs, KANs abandon explicit tensor-product or radial constructions in favor of factorized, additive representations, and instead rely on hidden layers and composition to approximate multivariate interactions. Case 3 demonstrates that once hidden layers are introduced, depth induces both multiplicative growth of polynomial degree and nonlinear coupling of coefficients, fundamentally departing from classical kernel structure. Together, these mechanisms explain the expressive power of deep KANs as well as the associated challenges in optimization, conditioning, and parameter identifiability, motivating kernel-inspired regularization and stability-aware training strategies.
@@ 4 Bridging KANs and MLPs
$\mathbf{z}^{(0)}$ $=\mathbf{x},$ (8)
where
  • $n_{\ell}$ is the width (number of neurons) of layer $\ell$,
@@
Formal Equivalence. Wang et al. [55] establish a precise bidirectional correspondence between multilayer perceptrons (MLPs) and Kolmogorov–Arnold Networks (KANs), showing that the two architectures are formally equivalent under mild structural assumptions.
Theorem 5 (MLP $\Rightarrow$ KAN, [55])
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain, and suppose that a function $f:\mathbb{R}^{d}\to\mathbb{R}$ can be represented by an MLP of width $W$, depth $L$, and activation function
$\sigma_{k}(x)=\max(0,x)^{k},\qquad k\geq 1.$
Then there exists a KAN of width $W$, depth at most $2L$, grid size $G=2$, and $k$-th order B-spline basis functions such that
$f(\mathbf{x})=g(\mathbf{x}),\qquad\forall\mathbf{x}\in\Omega.$
Concretely, an MLP representation of the form
$f(\mathbf{x})=\sum_{i=1}^{W}\alpha_{i}\,\sigma_{k}\!\Bigl(\sum_{j=1}^{d}w_{ij}x_{j}+b_{i}\Bigr)$
admits an equivalent KAN parameterization
$f(\mathbf{x})=\sum_{i=1}^{W}\Phi_{i}\!\biggl(\sum_{j=1}^{d}\varphi_{ij}(x_{j})\biggr),$
where each univariate map $\varphi_{ij}:\mathbb{R}\to\mathbb{R}$ is a learned B-spline that approximates the composite nonlinearity $\sigma_{k}(w_{ij}x_{j}+b_{i})$. From this perspective, KANs replace fixed pointwise activations in MLPs by learnable one-dimensional functions along each edge, without increasing the architectural width.
The reverse direction holds under mild restrictions on the univariate components. Non-polynomial nonlinearities (such as SiLU) cannot be captured exactly by $\mathrm{ReLU}^{k}$ networks, but if a KAN employs only polynomial-type univariate functions—such as B-splines of order kk—then an equivalent MLP representation exists.
Theorem 6 (KAN $\Rightarrow$ MLP, [55])
Suppose that a function $f:[0,1]^{d}\to\mathbb{R}$ is represented by a KAN of width $W$, depth $L$, grid size $G$, and $k$-th order B-spline univariate functions, and assume that no non-polynomial activations are present. Then there exists an MLP with activation $\sigma_{k}(x)=\max(0,x)^{k}$, depth at most $2L$, and width $(G+2k+1)W^{2}$ that represents $f$.
The induced MLP has parameter complexity $\mathcal{O}(G^{2}W^{4}L)$, whereas the original KAN requires only $\mathcal{O}(GW^{2}L)$ parameters. This gap highlights the superior parameter efficiency of KANs, particularly for fine spline grids or higher-order bases. As a consequence, approximation and convergence results developed for $\mathrm{ReLU}^{k}$ MLPs can be transferred directly to KANs through this formal equivalence (see Section 9).
@@
4.1 How KANs Extend the MLP Landscape
@@ 5.1 B-spline
Subfigures 2(a) and 2(b) visualize cubic (k=3k{=}3) B-spline bases on non–extended and extended grids, respectively. Without extension, boundary bases are truncated and lose symmetry; with extension, added knots outside the domain restore full polynomial support (gray bands). To compare expressivity, subfigures 2(c) and 2(d) synthesize univariate maps
@@
Spectral bias and NTK conditioning. Spline grids control frequency reach: finer grids reduce spectral bias but may worsen conditioning [55, 41]. Initialization also matters—Rigas et al. [283] [^15] https://github.com/srigas/KAN_Initialization_Schemes show that Glorot-style and power-law initializations yield more stable NTKs and faster early convergence. (See Secs. 9.2 and 9.3.)
@@ 5.3 ReLU
Figure 4: (a) Step-by-step construction of the normalized ReLU-based local basis function from its constituent $\operatorname{ReLU}(e_{i}-x)$ and $\operatorname{ReLU}(x-s_{i})$ terms. (b) Complete set of normalized ReLU–KAN basis functions $R_{i}(x)$ with supports $s_{i},e_{i}] [$129].